OpenStudy (anonymous):

assume that a perfessional baseball team has 11 ptchers,5 infielders, and 10 other players. if 3 players names are selected at random, determine the probl that all 3 are infielders.

OpenStudy (anonymous):

5 choose 3 divided by 21 choose 3

OpenStudy (anonymous):

so 3?

OpenStudy (anonymous):

is it 5/26?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Satellite already gave you the answer.

OpenStudy (anonymous):

is it 3?

OpenStudy (anonymous):

whats does 5 choose 3 devided by 21 choose 3 mean its only suppose to be 1 number rite

OpenStudy (anonymous):

Choose is the name we sometimes give to the binomial coefficient function: \[\text{n choose k} = {n\choose k} = \frac{n!}{k!(n-k)!}\]

OpenStudy (anonymous):

5!/3(21-3)!

OpenStudy (anonymous):

Not quite

OpenStudy (anonymous):

5 choose 3 = \(\frac{5!}{3!(5-3)!}\) 21 choose 3 = \(\frac{21!}{3!(21-3)!}\) Then divide those two values to find \[\frac{\text{5 choose 3}}{\text{21 choose 3}}\]

OpenStudy (anonymous):

18-7

OpenStudy (anonymous):

no

OpenStudy (anonymous):

because when u do 21/3 it equals 7

OpenStudy (anonymous):

Are you being taught to do this on your calculator using the Cr function?

OpenStudy (anonymous):

no we don't use our calculators for it

OpenStudy (anonymous):

Doing it by hand you have \[\frac{21!}{3!*18!} = \frac{21*20*19*18!}{3! * 18!} = \frac{21*20*19}{3*2} \]\[= 7*10*19 = 70*19 = 700 + 630= 1330\]

OpenStudy (anonymous):

But that will make for a very long evening.

OpenStudy (anonymous):

can i butt in for a moment?

OpenStudy (anonymous):

it as 1/260

OpenStudy (anonymous):

If i want to know the number of ways to choose 3 people (items, whatever) out of a set of 21 I reason as follows: there are 21 choices for the first person, 20 for the next and then 19 for the last. by the counting principle there are then 21*20*19 ways to do this. But this counts (a,b,c) differently than (b,c,a) for example so I have over-counted by the number of ways I can permute 3 things. There are 3*2 = 6 such permutations. Thus the number of ways to choose 3 from a set of 21 is 21*20*19/3*2. Similarly the number of ways to choose 3 from a set of 5 is 5*4*3/3*2 and the number of ways to choose 4 out of 11 is 11*10*9*8/4*3*2 etc

OpenStudy (anonymous):

Hello, Your chance of picking one is 5/26, then, once you've picked him, there are only 4 left out of 25 and once you've picked that player, there's only 3 left out of 24 to choose from (like in elementary school picking teams) giving you: (5/26)*(4/25)*(3/24) = 1/260, like you said.